OpenStudy (anonymous):

What is the rate of change for the linear relationship modeled in the table? x y -1 10 1 9 3 8 5 7

OpenStudy (anonymous):

@SamBrowne10

OpenStudy (anonymous):

I'm here

OpenStudy (anonymous):

Do you know how to do this @SamBrowne10

OpenStudy (michele_laino):

for example, going from x=3 to x=5, the change in y is 7-8=-1, so the rate of change is: \[\Large r = \frac{{7 - 8}}{{5 - 3}} = ...\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Every time x goes up 2 Y would go down 1 so....oh someone else already answered

OpenStudy (anonymous):

so the answer is -1/2?

OpenStudy (anonymous):

Just so you know i'm not still in school my teacher just hasn't submitted my final grade so i'm resubmitting assignments to get a better grade.

OpenStudy (anonymous):

Thank you for your help though!